# How structured are the representations in transformer-based vision encoders? An analysis of multi-object representations in vision-language models

### v2

@@ Abstract
Forming and using symbol-like structured representations for reasoning has been considered essential for generalising over novel inputs. The primary tool that allows generalisation outside training data distribution is the ability to abstract away irrelevant information into a compact form relevant to the task. An extreme form of such abstract representations is symbols. Humans make use of symbols to bind information while abstracting away irrelevant parts to utilise the information consistently and meaningfully. This work estimates the state of such structured representations in vision encoders. Specifically, we evaluate image encoders in large vision-language pre-trained models to address the question of which desirable properties their representations lack by applying the criteria of symbolic structured reasoning described for LLMs to the image models. We test the representation space of image encoders like VIT, BLIP, CLIP, and FLAVA to characterise the distribution of the object representations in these models. In particular, we create decoding tasks using multi-object scenes from the COCO dataset, relating the token space to its input content for various objects in the scene. We use these tasks to characterise the network’s token and layer-wise information modelling. Our analysis highlights that the CLS token, used for the downstream task, only focuses on a few objects necessary for the trained downstream task. Still, other individual objects are well-modelled separately by the tokens in the network originating from those objects. We further observed a widespread distribution of scene information. This demonstrates that information is far more entangled in tokens than optimal for them to represent objects similar to symbols. Given these symbolic properties, we show the network dynamics that cause failure modes of these models on basic downstream tasks in a multi-object scene.
@@
Abstraction Symbols Vision-Language Models Vision Encoder
@@ 1 Introduction
This study investigates whether symbol-like structured representations may emerge in transformer-based vision models. In vision models, some studies have evaluated their representations, looking for aspects of human-like symbolic cognitive processes [19, 21]. An advantage of searching for symbolic representations in vision models is that the input information is distributed in pixels, allowing explicit binding dynamics to be observed. Binding dynamics refers to the information about entities (objects, living things, see Fig 1.C) being bound together into smaller representational units, describing specific useful aspects of that entity [9]. Further, in vision models, the intermediate representations of objects are natural candidates to be formed into symbols, allowing one to formulate hypotheses about the nature of the representations. Building upon the notion of symbols and binding described in previous studies [9, 19, 21], we propose three levels of properties for an image encoder representation to be more symbolic: 1) The model should be able to aggregate information specific to objects in the image into smaller units; 2) The model should represent various logical components of the input image in separate computational units, i.e., there should be object-wise information disentanglement; 3) A downstream processing network should be able to utilise the composition of these units for the required task, e.g., scene modelling for captioning.
@@
The first two of these proposed properties of being symbol-like promote better out-of-distribution generalisation by themselves. 1) Better binding of object information into a discrete set of units allows useful properties of the representation to be fully available for novel scenarios in solving downstream task. 2) Disentanglement of these units of representation makes sure that the presence of one object does not affect the representation of others. This is important for extreme cases where the objects are present out of the scene statistics of the trained distribution and these spurious correlations in representations of object presence may cause failure for the downstream task.
The compositional abilities of these Vision-Language Models (VLMs) have already been tested and shown to be inadequate [28]. However, there is some evidence that while doing downstream tasks, primitive concepts emerge in model representations and are compositionally employed by the downstream network [29]. Transformers have already been shown to have global and local information aggregation within a single layer facilitated by the attention mechanism [24]. Also, due to the retention of the token space across layers, they could have enough computational units to maintain object information separately while binding them locally. Hence, there is a need to estimate the extent to which the first two properties of symbolic representations listed above are implemented in vision transformer models.
@@
For this purpose, we analyse the spatial representations in the vision transformers’ token space. Our primary questions testing for properties 1 and 2 are the following: Do transformers represent and maintain object-wise representations? Are these representations disentangled, i.e. does a particular set of tokens only represent a particular object? We utilise the COCO dataset[17] with its instance object masks to characterise the token representations in relation to their input patch information. We set up decoding experiments in a two-object setting (in a multi-object scene) to determine how the encoders manage the representations of the two objects. We try encoder of three VLMs (CLIP [23], BLIP[16] and FLAVA[25]) and compare them with the larger versions (CLIP-L, BLIP-L) and also check the representations against VIT trained for image classification and CLIP (Resnet X4) with a CNN backbone.
We generally observe that object-specific areas hold the most discriminative features about the objects until the last layers. The token representations are not disentangled since they can decode other objects in the scene with accuracy far above random guesses. Across networks, we see similar patterns of decoding accuracies in transformer-based VLMs. We observe the VLMs trained on objectives requiring the modelling of multiple objects have better symbol-like differentiated object representations than VIT encoders trained for image classification. On the other hand, all VIT-based encoders have less differentiated and disentangled object representation in CNN-based CLIP (Resnet X4). Our results suggest a preference for modelling objects relevant to the downstream task (e.g., captioning for BLIP). This effect is more pronounced for the CLS token directly optimised for the task; other object-specific tokens still retain more high-level information about most objects. We analyse and discuss the implications of representation loss in CLS for failure in out-of-distribution task scenarios.
2 Related Works
The overall problem with generalisation in DL methods and the need for explicit inductive biases that allow for discrete yet flexible information binding has been comprehensively discussed [9]. Discrete symbolic representations are considered a prerequisite for robust compositionality and reasoning [21] and new studies propose that such discrete information binding may originate from training existing models on large datasets [21, 19]. This issue carries over to the field of meta-learning with its aim to introduce and compositionally utilise modularity in neural networks to improve generalisation [13].
@@ 2 Related Works
Concept bottleneck models are a series of models that try to explicitly learn human interpretable intermediate outputs to be composed into final output labels [7, 27]. This interpretability usually comes at the cost of downstream task performance. [11] introduced an extra loss in the intermediate layer on the networks for units to align to the interpretable concepts while preserving task performance. Many other works do not bind representations to any explicit intermediate concept but still have explicit discrete bottlenecks in their networks [20, 18]. It has been shown that these networks learn better representations which further help generalisation on downstream tasks [20, 18, 26].
2.1 Evaluating model’s reasoning capabilities
@@
Many works have proposed benchmarks, and others have tried to analyse the existing networks to estimate the reasoning capabilities of the ANN models. [3] proposed a visual question-answering benchmark to evaluate the reasoning skills on images, with open-ended and free-form questions and expected solutions in free-form natural language. Visual genome [12] provided content-rich images with explicit annotation of object position and relationships, promoting models that exploit such information. Datasets such as CLEVR [10] were designed to correct for biases the models would utilise in the existing benchmarks to perform well without explicit reasoning. The ARO (Attribute, Relation, Order) dataset [28] has been proposed to test the correct binding of information when compositionally representing multi-object scenarios. Particularly in VLMs such as BLIP and CLIP, they find that the model outputs do not bind the compositional properties well and attribute the wrong order and features to objects when describing them.
@@ 2.2 Interpreting model’s representations
In terms of analysis of model representations for concepts, early work [2] showed the layer-wise progression of decoding accuracy for concepts in CNNs. [6] reveals the inner workings of transformers by showing how trained transformers can implement convolutions and, in the initial layers, form grid-like local attention patterns like a convolution filter. [24] shows that vision transformers differ from CNNs because of their ability to encode both local and global information in the initial layer. Instead, in CNNs, there is a multi-scale feature representation going from lower-scale local information captured in the initial layers to higher-scale global information captured in the higher layers [4].
@@
In contrast to the generic visualisation tools that look into the model’s functioning, many studies inspect their representations, evaluating a specific computational functionality. [14] finds modular subnetworks in trained ANN models functionally responsible for separate tasks. Multiple studies further look into the notion of concepts in trained vision-language model representations. One study [29] designed a test to check if primitive concepts emerge in the network’s representations, which are used compositionally for downstream tasks. [19] defines tests for concepts in visual representations according to Fodor’s criteria [8] and tests these criteria using a controlled synthetic dataset. In our work, we build on the intuitions from these works while trying to formulate our hypothesis about symbolic representations in a limited scope but testing it with images from natural scenes.
@@ 3 Methods
Figure 1: A. Explanation of how the token representations are obtained. We analyse four kinds of tokens in this study 1) CLS token: Is common to both objects in the image; Object-specific 2) Avg_obj: obtained by averaging the token representations of the object-masked tokens as shown in the figure. 3) Random_obj (Object token): Rather than averaging, we sample one of the tokens from the masked token space of the object 4) Random : Obtained by sampling any random token from the token space other than the CLS token B. Describes the experimental setup in which we perform decoding in paired object tasks; each object-specific representation decodes 1) the object itself, 2) the other object in the image, and 3) the combination of both objects. C. Shows a sample paired object decoding task; given an image, the task is to decode if it contains object1 (cat/dog), object2 (chair/couch) or a combination of both.
@@
We design probes and use similarity measures to evaluate the network’s representational structure. In the following section, we describe the details of the experimental setup, data and the networks analysed.
Experimental Setup: We decode the combination of objects in the image from a single token or an average of tokens obtained from various parts of the image (see Fig. 1.A). The tokens we are interested in include the 1) CLS token (the token used and trained for the downstream task), 2) Average token (avg_obj): the average of token representation obtained from the object, 3) Random object (random_obj) token: a single token randomly sampled from the object 4) Random token (Random): a randomly sampled token from the image that served as a baseline. The token representation is obtained at the output of each layer. In this process, to identify the tokens originating from an object, we scale the segmentation mask of the object to the size of the token space. The probing task is designed as a classification task with the following settings: we use the tokens originating from 1) Primary object, 2) Secondary object and use it to decode 1) Primary object category, 2) Secondary object category, 3) Combination of primary and secondary object categories (see Fig. 1.B). We train and evaluate the probes with a train/val/test setup with 80/10/10 percent data splits. The reported accuracies are all final test set accuracies. All trained probes are linear and use the Scikit-learn[22]’s perceptron implementation, with its default parameters. To check the baseline performance of these representations to decode objects in a complex (20 classes) task, we train layer-wise global object classification probes for each layer of the networks. These probes utilise first 40000 images from the MSCOCO train set for training the probes and the 5000 images in the validation set for testing. To make the linear probe with less learning capacity, we use only the top 20 frequent object categories in the 40000 training images in the probing task.

### v3

# Interpreting the structure of multi-object representations in vision encoders

@@ Abstract
In this work, we interpret the representations of multi-object scenes in vision encoders through the lens of structured representations. Structured representations allow modeling of individual objects distinctly and their flexible use based on the task context for both scene-level and object-specific tasks. These capabilities play a central role in human reasoning and generalization, allowing us to abstract away irrelevant details and focus on relevant information in a compact and usable form. We define structured representations as those that adhere to two specific properties: binding specific object information into discrete representation units and segregating object representations into separate sets of tokens to minimize cross-object entanglement. Based on these properties, we evaluated and compared image encoders pre-trained on classification (ViT), large vision-language models (CLIP, BLIP, FLAVA), and self-supervised methods (DINO, DINOv2). We examine the token representations by creating object-decoding tasks that measure the ability of specific tokens to capture individual objects in multi-object scenes from the COCO dataset. This analysis provides insights into how object-wise representations are distributed across tokens and layers within these vision encoders. Our findings highlight significant differences in the representation of objects depending on their relevance to the pre-training objective, with this effect particularly pronounced in the CLS token (often used for downstream tasks). Meanwhile, networks and layers that exhibit more structured representations retain better information about individual objects. To guide practical applications, we propose formal measures to quantify the two properties of structured representations, aiding in selecting and adapting vision encoders for downstream tasks. Overall, we aim to advance the understanding of object-wise structured representations in vision encoders, thus enhancing their transparency and interpretability. By clarifying how these models bind and segregate object-level information, we enable better-informed decisions for optimal downstream task adaptation, ultimately aligning their behaviour more closely with human reasoning.
@@
Vision Encoders Interpretability Structured Representations Transparency Token Analysis Explainable AI Transformers.
@@ 1 Introduction
Vision encoders are foundation models for complex visual inputs in AI, including applications where decisions must be transparent and justifiable, such as healthcare, autonomous vehicles, etc. However, the opacity of these models limits their trustworthiness and applicability in real-world scenarios [23]. Recognizing the necessity of increasing direct transparency of vision encoders that cannot be guaranteed by post-hoc methods [33], it becomes essential to focus on the structural representations within the vision encoders themselves. Vision encoders are trained under various objective functions to learn suitable representations of visual inputs, and a good representation is the one that contains the correct details for its downstream task [5]. Humans flexibly use structured mental representations for reasoning, where the exact aspect of the representation is generated according to the task context [31, 17]. For instance, in a multi-object scene, we usually focus on only the main objects and infer their relationship to make inferences about the overall situation. However, we also have the ability to focus on individual objects in the scene if the task requires such richer representation abstracted from the scene; e.g., questions about the colour of a chair should be answered irrespective of what scene or surrounding one sees the chair.
In computer vision, especially for practical applications like robotics, vision encoders trained on large datasets are also expected to be useful for both the scene and the object-level encoding [34]. To understand to what extent the representations of pre-trained vision encoders allow for such ability, this study investigates if structured object-wise representations exist in the token space of vision encoders. As previous works [12, 25, 28] discuss how such abilities are dependent on having structured representations in a representation system, we build upon this notion of symbols and binding of representation described in these studies to propose two properties specific for an image encoder’s representation to be more structured in token space:
  1. 1.
    ($M_{1}$) The model should be able to bind information specific to objects in the image into specific representation units, i.e. a fewer number of tokens can represent the object better than its input token representation;
  2. 2.
    ($M_{2}$) The model segregates representation of various objects of the input image in separate sets of tokens, i.e., there should be object-wise information disentanglement in the token space.
@@
We expected these two properties for structured representation to promote better generalisation for the following reasons. First, binding object information into discrete units allows beneficial representation properties to be fully available for novel scenarios in solving downstream tasks. Second, disentanglement of these representation units ensures that one object’s presence in the scene does not affect others’ representation. This is important for extreme cases where the objects are present out-of-the-scene statistics of the trained distribution. Such spurious correlations in representations of object presence may cause failure for the downstream task.
@@
As we analyse the spatial representations in the vision transformers’ token space, our primary questions testing the above properties are the following: Do transformers represent and maintain object-wise representations? Are these representations disentangled, i.e. does a particular set of tokens only represent a particular object? We use the COCO dataset [22] with its instance object masks to characterise the token representations in relation to their input patch information. We set up decoding experiments in a two-object setting (at a time, within a multi-object scene) to determine how the encoders manage the representations of the two objects. We test encoders of three VLMs (CLIP [30], BLIP [21] and FLAVA [35]) and compare them with the larger versions (CLIP-L, BLIP-L), and also check the representations against a CLIP (Resnet X4) with a CNN backbone. We further compare against other pre-training objectives such as VIT (trained for image classification), DINO [8] and DINOv2 [27] (encoders pre-trained by image self-supervision).
To systematically evaluate these two properties, we design complementary tasks: (1) a paired-object decoding setup focusing on two objects at a time from the scene, and (2) a broader 20-object decoding task testing generalization to unrelated categories. In this way, we validate the inferences made by the original task and illustrate the usefulness of our proposed measure to predict the general object encoding ability of the structured representations.
In the literature, some experimental results were given alongside the release of the original models, which already provide partial inferences about their representations. For example, in CLIP [30], an experiment shows how learning linear decoder on the CLS token representation performs better than their proposed zero-shot classification. In DINO [8], we can see how different attention heads from the CLS token attend to prominent objects in the scene. While in DINOv2 [27], each token from the last layer can be classified using a linear layer to perform semantic segmentation to some satisfactory detail, showing the object-wise information expressed in the token space of the last layer.
In our analysis, motivated by these prior indications, we conduct a more systematic investigation into how object-specific representations arise across various encoders. We also test their generality, aiming to provide a more comprehensive understanding of the structure of information expressed in the token space of vision transformers that explains such observations. Our results show that the CLS tokens best models image-level information but, as an effect, prefer to represent the main objects in the scene. Still, object-specific areas in the tokens of the higher layers model hold the most discriminating features about the individual objects. Yet, the object-wise token representations are not disentangled since they can decode other objects in the scene with accuracy far above random guesses. We observe that the VLMs trained on objectives requiring the modeling of multiple objects have better multi-object representations in CLS token than VIT encoders trained for image classification. On the other hand, all VIT-based encoders have less segregated and more entangled object representations than CNN-based CLIP (Resnet X4). We highlight the CLS token’s inability to model background objects in all models (except CLIP-L) as a possible failure mode when using this token for downstream tasks on multi-object images. Moreover, we show that our two measures of structured representations in the token space of the networks correlate highly with the retention of object-specific information for (less important) background objects. Hence, it illustrates the usefulness of these structured representations in representing multi-object scenarios in the token space and provides a measure to make decisions about appropriate layers to train decoders or adaptors for an appropriate downstream task.
Our main contributions are:
  1. 1.
    We identify two properties to evaluate structured representations in the token space of image encoders. We design experiments and create a relevant dataset using COCO’s multi-object images to evaluate these properties and propose a measure related to each property.
  2. 2.
    Our method lets us understand the nature of information encoding through the lens of structured representations in vision encoders in the token space (about the encoded objects) at different encoder layers.
  3. 3.
    The results from this analysis led us to identify a failure in generalising over multiple objects in the scene. Further, our proposed measures of structured representation ($M_{1},M_{2}$) help against the problem by indicating which tokes/layers and the kind of models can be used to obtain the best representations for the (less represented) background objects.
2 Related Work
@@
Our investigation into structured representations in vision encoders relates to a greater problem of any model’s ability to generalise across different scenarios. This overall problem with generalisation in deep learning (DL) methods has been comprehensively discussed [12] in light of the need for explicit inductive biases that allow for discrete yet flexible information binding. Discrete symbolic representations are considered a prerequisite for robust compositionality and reasoning [28], and new studies propose that such discrete information binding may originate from training existing models on large datasets [28, 25]. Beyond discrete symbolic approaches, this issue carries over to the field of meta-learning with its aim to introduce and compositionally utilise modularity in neural networks to improve generalisation [17]. Concept bottleneck models are a series of models that try to explicitly learn human-interpretable intermediate outputs to be composed into final output labels [10, 38]. However, this interpretability usually comes at the cost of downstream task performance. [15] introduced an extra loss in the intermediate layer on the networks for units to align to the interpretable concepts while preserving task performance. Many other works do not bind representations to any explicit intermediate concept but still have explicit discrete bottlenecks in their networks [26, 24]. It has been shown that these networks learn better representations, which further help generalisation on downstream tasks [26, 24, 36].
2.1 Evaluating a trained model’s capabilities
@@
In the case of evaluation of pre-existing networks, many works do not try to estimate the mechanisms or representation structure that lead to particular performance of networks on tasks but rather propose benchmarks to estimate the reasoning capabilities of the ANN models. [3] proposed a visual question-answering benchmark to evaluate the reasoning skills on images, with open-ended and free-form questions and expected solutions in free-form natural language. Visual genome [16] provided content-rich images with explicit annotation of object position and relationships, promoting models that exploit such information. Datasets such as CLEVR [14] were designed to correct for biases the models would utilise in the existing benchmarks to perform well without explicit reasoning. The ARO (Attribute, Relation, Order) dataset [39] has been proposed to test the correct binding of information when compositionally representing multi-object scenarios. Particularly in VLMs such as BLIP and CLIP, they find that the model outputs do not bind the compositional properties well and attribute the wrong order and features to objects when describing them. Though these benchmarks reveal model shortcomings at the final output, they do not clarify how (or if) the internal representations disentangle multiple objects.
@@ 2.2 Interpreting model’s representations
In terms of analysis of model representations for concepts, early work by [2] showed the layer-wise progression of decoding accuracy for concepts in CNNs. [9] reveals the inner workings of transformers by showing how trained transformers can implement convolutions and, in the initial layers, form grid-like local attention patterns like a convolution filter. [32] shows that vision transformers differ from CNNs because of their ability to encode both local and global information in the initial layer. In contrast, CNNs exhibit a multi-scale feature representation hierarchy going from lower-scale local information captured in the initial layers to higher-scale global information captured in the higher layers [6].
@@
In contrast to these generic visualisation tools that look into the model’s functioning or attention-based approaches, other studies inspect model representations, evaluating a specific computational functionality. [18] finds modular sub-networks in trained ANN models functionally responsible for separate tasks. Multiple studies further look into the notion of concepts in trained vision-language model representations. A study by [40] designed a test to check if primitive concepts emerge in the network’s representations, which are used compositionally for downstream tasks. [25] defines tests for concepts in visual representations according to Fodor’s criteria [11] and tests these criteria using a controlled synthetic dataset. Building on these insights, our work formulates explicit hypothesis about structured representations regarding object-wise binding and disentanglement in a well-defined scope of layer-wise distribution of representations in visual encoders. We explore natural scenes rather than heavily controlled settings, thus complementing the more synthetic or attention-centered studies above.
@@ 3 Methods
Figure 1: A. Explanation of how the token representations are obtained. We analyse four kinds of tokens in this study: 1) CLS token: The special token usually used in models for downstream tasks; 2) Avg_obj(Object-specific token): obtained by averaging the token representations of the object-masked tokens as shown in the figure. 3) Random_obj (Object-specific token): Rather than averaging, we sample one of the tokens from the masked token space of the object 4) Random: Obtained by sampling any random token from the token space other than the CLS token B. Describes the experimental setup in which we perform decoding in paired object tasks; each object-specific representation decodes 1) the object itself, 2) the other object in the image, and 3) the combination of both objects. C. Shows a sample paired object decoding task; given an image, the task is to decode if it contains object1 (cat/dog), object2 (chair/couch) or a combination of both.
@@
We train probes and use similarity measures to evaluate the network’s representational structure. In the following section, we formally define measures related to our proposed properties of structured representations and describe the details of the experimental setup, data, and the networks analysed.
3.1 Formal Definition
Our method quantifies the information structure in layers of an image encoder by measuring two properties. Given an image $I$, containing at least two objects, denoted as a primary object $O_{p}$ and a secondary object $O_{s}$, we denote the tokens originating from the primary object as $T_{p}$ and those originating from secondary objects as $T_{s}$. Across a test set, the accuracy of a probe trained and evaluated using representation from $T_{p}$ to decode object $O_{s}$ is given by $A_{ps}$. Similarly, $A_{pp}$,$A_{sp}$ and $A_{ss}$ refer to the other token-object combinations.
In this study, for the first property (binding), we measure the ability of a localised region to decode the individual object. More precisely, we focus on tokens $T_{s}$ ’s ability to decode object $O_{s}$ linearly. Hence, we measure this over a set of images by $A_{ss}$. Thus,
$M_{1}=A_{ss}$
The second property (segregation) estimates the mixing of information between the tokens from other objects. We estimate if representations from tokens $T_{s}$ can decode object $O_{p}$. Over a set of images, we use the decoding accuracy $A_{sp}$ to estimate this entanglement between token representations, which is the opposite of segregation. For better comparison among layers, we scale this value by the decoding accuracy of that object for from own token representation, i.e. $A_{pp}$. Hence, the final measure for entanglement of information is given by $\frac{A_{sp}}{A_{pp}}$.
$M_{2}=\frac{A_{sp}}{A_{pp}}$
3.2 Experimental Setup: Paired Object Task
The main experiment for assessing how token space encodes multi-object information is the paired object task. The objective of this experimental setup is to have images that contain only one of the two primary objects and one of the four secondary objects. The basic inference we make is that given a particular token representation of the image from the encoder, can we reliably infer both the primary and the secondary object?
We try to infer the multi-object relations in a pairwise manner as the objects vary in location, size and occur in different scene context. We select a total of 6 object sets that help generalise over different object types. The particular importance of the objects is selected by their relative importance to the scene, which is further formalised in later experiments by quantifying their mention in the COCO captions for the particular image. We decode the combination of objects in the image from a single token or an average of tokens obtained from various parts of the image (see Fig. 1.A). The tokens we are interested in include are:
  1. 1.
    CLS token: the token used and trained for the encoder’s downstream task,
  2. 2.
    Average token (avg_obj): the average of token representation obtained from the object,
  3. 3.
    Random object token (random_obj): a single token randomly sampled from the object,
  4. 4.
    Random token (Random): a randomly sampled token from the image that served as a baseline.
The token representation is obtained at the output of each layer. To identify the tokens originating from an object, we scale the segmentation mask of the object to the size of the token space.
The paired-object probing task is designed as a classification task with the following settings: we use the tokens originating from 1) Primary object ($T_{p}$), 2) Secondary object ($T_{s}$) and use it to decode 1) Primary object $(O_{p}$) category, 2) Secondary object ($O_{s}$) category, 3) Combination of primary and secondary object categories (see Fig. 1.B). We train and evaluate the probes with a train/val/test setup with 80/10/10 percent data splits. The reported accuracies are all final test set accuracies. All probes are linear and use the Scikit-learn [29]’s perceptron implementation, with its default parameters.
3.3 Experimental Setup: Global Probe
The global decoding task is proposed to test the generalisation of the numbers obtained by the paired-object task on completely different categories of objects. It also allows us to verify using a more complex multi-class (20 class) probe that the trends in the paired object task are not a result of a simple 2 class or 4 class decoding (as in the paired-object task). Finally, in this analysis, we formalise another notion of object importance in the test set data split between objects mentioned in the caption as ’in the caption’ and objects not mentioned in the caption as ’not in the caption’. This further helps test the difference between main and secondary objects apart from the category level classification as primary and secondary objects used in the paired object task.
In this task, object-specific probes are trained, i.e. using Random_Obj and Avg_obj tokens to decode the object category from which those tokens are taken. These probes utilise the first 40000 images from the MSCOCO train set to train a probe for the layer and each token type. The 5000 images in the validation set are used for testing. As we are testing for generalisation, we exclude the object categories used in the six object sets used to calculate our structured representation measures given in Table 1 and randomly choose 20 object categories (given in Global section of Table 1) in the COCO dataset for this evaluation.
3.4 Dataset